# Out of Style: RAG’s Fragility to Linguistic Variation

### v1

@@ 5.2 Generation Analysis
As illustrated in Table 3, across all datasets, we observe a noticeable overall degradation in performance when queries are rewritten to become less formal, more polite, less readable, or have grammatical errors. Across all datasets, linguistic dimensions, and experimental settings, we found average drops of 16.52% (AM), 41.15% (EM), and 19.60% (F1). The PopQA dataset shows the highest sensitivity to all linguistic variations, with an average performance drop of 18.64% on AM scores. Particularly notable were the effects of reduced readability (18.22% degradation) and round-trip translation (33.86% degradation). These findings suggest that while the RAG systems perform well on standard NLP benchmarks with structured queries, they remain vulnerable to common linguistic variations. The full experiment results are shown in Appendix D.
@@
When queries are rephrased to be more polite, we find that the AM scores remain relatively similar to those of the original, with less than 10% change for all datasets. However, there are significant drops in exact match (EM) scores. Specifically, we observe 44.57% and 18.32% drops in EM scores for queries from the Natural Questions and PopQA datasets, respectively. Through manual checks, we find that LLMs tend to generate more complete or formal responses in polite query formulations, which results in lower EM rates but similar performance on AM.
@@ 5.4 Exploring the Robustness of Advanced RAG Systems
Many modern RAG systems include more components than simply retrieval and generation which aim to make them more useful for users (Gao et al., 2022). In our experiments, we broadly explore whether the addition of two advanced components—Query Expansion and Reranking—improves RAG robustness. Detailed results of the experiment can be found in Appendices E and F.
@@ 5.4.2 Reranker
The retriever must be efficient for large document collections containing millions of entries, although it may sometimes retrieve irrelevant candidates. To address this, we incorporate a Cross-Encoder-based re-ranker to significantly enhance the quality of final answers. Specifically, we employ the MS MARCO Cross-Encoders developed by Reimers & Gurevych (2019) to re-rank passages retrieved by ModernBERT and Contriever. As illustrated in Figure 3, re-ranking substantially improves retrieval performance, particularly for rewritten queries, achieving an average improvement of 16.56% compared to only 7.40% for original queries. In contrast, generation results show a modest improvement of 1.83% for original queries and a more substantial improvement of 8.50% for rewritten queries. These findings suggest that the effectiveness of re-ranking is especially pronounced when handling rewritten queries and highlight the importance of improving the robustness of retrieval systems.
@@ A.1 Linguistic Variations
Each rewrite for each prompt must adhere to these thresholds in order to develop the final dataset we use. In addition to this, we want to ensure that the rewritten queries are semantically similar to the original query. Therefore, in the query-rewriting process, we add an additional constraint where the sentence similarity between the original and rewritten queries must be greater than 0.7, as assessed by the MPNet-v2 model Song et al. (2020) in the SentenceBert Library Reimers & Gurevych (2019).
@@
FRES=206.8351.015(total wordstotal sentences)84.6(total syllablestotal $\text{FRES}=206.835-1.015\left(\frac{\text{total words}}{\text{total sentences}}\right)-84.6\left(\frac{\text{total syllables}}{\text{total words}}\right)$ (1)
@@ A.2 Semantic Preservation of Linguistic Variations
To ensure our linguistic rewrites preserve semantic meaning, we conducted a qualitative evaluation where two researchers annotated 20 randomly sampled query rewrites per linguistic property for the EntityQuestions dataset, totaling 100 queries examined by both annotators. Our primary assessment criterion was whether the rewritten question remained semantically similar enough to warrant the same answer as the original query.
For typos and back-translation variations specifically, we established a threshold of maximum 2 character edits on entities in the questions to determine if a query remained semantically equivalent to its original form. This threshold helps maintain a balance between introducing realistic linguistic variations and preserving essential meaning.
Our analysis revealed high semantic preservation rates across all linguistic dimensions:
Linguistic Dimension Semantic Preservation (%) Inter-annotator Agreement (%)
@@
Round-trip Translation 95.0 100.0
Table 4: Percentage of queries that maintain semantic similarity after linguistic rewrites and inter-annotator agreement rates.
These results demonstrate the robustness of our rewriting methodology in preserving semantic content while introducing natural linguistic variations. The high inter-annotator agreement further validates the reliability of our assessment approach. This semantic preservation is critical for our experimental design, as it ensures that performance differences observed in RAG systems can be attributed to linguistic variations rather than semantic drift.
@@ B.1 LLMs Generation Hyperparameter Setting
Appendix C Full Retrieval Experiment Results
@@
C.1 Scaling Number of documents
@@
Appendix D Full RAG Experiment Results
@@
Appendix E Full Query Expansion Experiment Results
@@
Appendix F Full Reranking Experiment Results
Dataset Linguistic Dimension R@5 R@10 R@20 R@100
@@ Appendix F Full Reranking Experiment Results
Appendix G Rewriting Prompts
G.1 Formality
@@ G.3 Politeness
Appendix H LLMs Prompts
H.1 Few-shot Prompts
@@
Appendix I Rewriting Examples

### v2

@@ 5.2 Generation Analysis
As illustrated in Table 3, across all datasets, we observe a noticeable overall degradation in performance when queries are rewritten to become less formal, more polite, less readable, or have grammatical errors. Across all datasets, linguistic dimensions, and experimental settings, we found average drops of 16.52% (AM), 41.15% (EM), and 19.60% (F1). The PopQA dataset shows the highest sensitivity to all linguistic variations, with an average performance drop of 18.64% on AM scores. Particularly notable were the effects of reduced readability (18.22% degradation) and round-trip translation (33.86% degradation). These findings suggest that while the RAG systems perform well on standard NLP benchmarks with structured queries, they remain vulnerable to common linguistic variations. The full experiment results are shown in Appendix F.
@@
When queries are rephrased to be more polite, we find that the AM scores remain relatively similar to those of the original, with less than 10% change for all datasets. However, there are significant drops in exact match (EM) scores. Specifically, we observe 44.57% and 18.32% drops in EM scores for queries from the Natural Questions and PopQA datasets, respectively. Although the rewritten queries preserve the original pragmatic intent, they introduce longer phrasing and auxiliary words that shift embeddings and sometimes nudge generators toward more verbose outputs. As a result, they still cause slight regressions in the RAG.
@@ 5.3 Retrieval Method and LLM Scale Influence
Human annotator evaluation aligns with our automated evaluation metrics
To validate our automated metrics and control for potential confounds such as length or verbosity, two independent annotators evaluated 250 samples across EntityQuestions (100 samples) and NaturalQuestions (150 samples). For typos and RTT variations, rewrites with more than 2 character edits to any entity were marked as not preserving semantic meaning. As shown in Table 5, semantic preservation rates ranged from 86.7% to 100% with high inter-annotator agreement (96.7%-100%), confirming that observed performance differences stem from linguistic variation rather than semantic drift. Full annotation protocols are provided in Section A.2.
@@ 5.4 Exploring the Robustness of Advanced RAG Systems
Many modern RAG systems include more components than simply retrieval and generation, which aim to make them more useful for users (Gao et al., 2022). In this section, we explore the possibility of using a simple query-expansion step to fix the vulnerability on linguistic variations and whether the addition of reranking improves RAG robustness. Detailed results of the experiment can be found in Appendices G and H.
@@ 5.4.2 Reranker
The retriever must be efficient for large document collections containing millions of entries, although it may sometimes retrieve irrelevant candidates. To address this, we incorporate a Cross-Encoder-based re-ranker to significantly enhance the quality of final answers. Specifically, we employ the MS MARCO Cross-Encoders developed by Reimers and Gurevych (2019) to re-rank passages retrieved by ModernBERT and Contriever. As illustrated in Figure 3, re-ranking substantially improves retrieval performance, particularly for rewritten queries, achieving an average improvement of 16.56% compared to only 7.40% for original queries. In contrast, generation results show a modest improvement of 1.83% for original queries and a more substantial improvement of 8.50% for rewritten queries. These findings suggest that the effectiveness of re-ranking is especially pronounced when handling rewritten queries and highlight the importance of improving the robustness of retrieval systems.
@@ 6 Conclusion
Limitations
While our choice of linguistic dimensions cover a broad spectrum of stylistic, pragmatic, and structural variations, other relevant factors such as dialect, idiomatic expressions, or domain-specific terminology could be explored in future work. Our grammatical variations (typos and round-trip translation) represent a subset of potential errors; other categories such as verb form or agreement errors Dahlmeier et al. (2013) may present greater challenges and warrant future investigation. We conducted query rewriting using two LLMs (GPT-4o-mini and Llama-3.1-70B-Instruct) and observed similar vulnerabilities; future studies may verify the generalizability of these findings using a broader range of rewriting methods. Additionally, while we explored widely used methods such as query expansion and reranking to test for mitigation strategies, more comprehensive approaches, including training models explicitly on diverse, linguistically varied data, remain important avenues for future research.
Acknowledgments
This research was supported in part by the National Science Foundation under grant 2230466 and in part by DSO National Laboratories.
@@ A.1 Linguistic Variations
Each rewrite for each prompt must adhere to these thresholds in order to develop the final dataset we use. In addition to this, we want to ensure that the rewritten queries are semantically similar to the original query. Therefore, in the query-rewriting process, we add an additional constraint where the sentence similarity between the original and rewritten queries must be greater than 0.7, as assessed by the MPNet-v2 model Song et al. (2020) in the SentenceBert Library Reimers and Gurevych (2019).
@@
$\begin{split}\mathrm{FRES}&=206.835-1.015\;\frac{\text{total words}}{\text{total sentences}}\\ &\quad-84.6\;\frac{\text{total syllables}}{\text{total words}}\end{split}$ (1)
@@ A.2 Semantic Preservation of Linguistic Variations
To ensure our linguistic rewrites preserve semantic meaning, we conducted a qualitative evaluation where two independent researchers annotated randomly sampled query rewrites per linguistic property from EntityQuestions (100 examples) and Natural Questions (150 examples) datasets, totaling 250 queries.
A.2.1 Annotation Guidelines
Annotators were provided with the following protocol to assess whether each rewritten query preserved the semantic meaning of the original:
  • Mark as semantically preserved if both queries would elicit identical factual answers
  • Mark as not preserved if the rewrite changes entities, relationships, or temporal/spatial constraints
  • Rule for Typos and RTT: If more than 2 characters are edited in any entity mention, mark as not preserved. This threshold helps maintain a balance between introducing realistic linguistic variations and preserving essential meaning.
  • Ignore stylistic differences that do not affect the core information need
Annotators worked independently and their judgments were compared to compute inter-annotator agreement.
Linguistic Variation Semantic Inter-annotator
Preservation (%) Agreement (%)
@@ A.2.1 Annotation Guidelines
RTT 95.0 100.0
(a) EntityQuestions
Linguistic Variation Semantic Inter-annotator
Preservation (%) Agreement (%)
Politeness 100.0 100.0
Formality 90.0 96.7
Readability 100.0 100.0
Typos 86.7 96.7
RTT 96.7 96.7
(b) Natural Questions
Table 5: Percentage of queries that preserve semantic meaning after linguistic rewrites, along with inter-annotator agreement rates across 250 annotated samples.
A.2.2 Results and Analysis
Table 5 shows high semantic preservation rates across all linguistic dimensions, ranging from 86.7% to 100%. The high inter-annotator agreement (96.7%-100%) further validates the reliability of our assessment approach. These results demonstrate the robustness of our rewriting methodology in preserving semantic content while introducing natural linguistic variations. Notably, even with our strict 2-character edit threshold for typos and RTT, we achieve high preservation rates (86.7%-96.7%), indicating that our generation process successfully balances realism with semantic fidelity. This semantic preservation is critical for our experimental design, as it ensures that performance differences observed in RAG systems can be attributed to linguistic variations rather than semantic drift.
@@ B.1 LLMs Generation Hyperparameter Setting
Appendix C Experiment Results on Llama 3.1 Rewriting Queries
We conducted supplementary experiments by sampling 500 queries from the PopQA dataset. We generated rewritten queries using Llama-3.1-70B-Instruct, employing the same rewriting criteria as previously described. Then, we performed RAG experiments using ModernBERT as the retriever and Qwen2.5-7B-Instruct as the generator. The results are summarized below:
Dimension Original Rewritten $\boldsymbol{\Delta}$
Readability 0.624 0.490 21.5%
Formality 0.826 0.746 9.7%
Politeness 0.728 0.626 14.0%
Table 6: Retrieval Performance (R@5)
Dimension Original Rewritten $\boldsymbol{\Delta}$
Readability 0.500 0.372 25.6%
Formality 0.776 0.684 11.9%
Politeness 0.604 0.520 13.9%
Table 7: RAG Generation Performance (Answer Match - AM Score)
In general, the RAG system remains notably sensitive to the linguistic variations introduced by LLaMA. Specifically, queries rewritten for reduced readability caused the most significant performance degradation, with 21.5% in retrieval accuracy and 25.6% in the AM score. Unlike variations introduced by GPT, the system exhibited greater sensitivity to changes in politeness compared to formality. We will present a comprehensive analysis and more detailed results in the revised paper.
Appendix D Data and Code Availability
Our code and rewritten query datasets will be released after the peer review stage under the CC BY 4.0 license. All existing datasets, models, and codes used in this work were employed consistently with their intended research purposes.
Appendix E Full Retrieval Experiment Results
@@
E.1 Scaling Number of documents
@@
Appendix F Full RAG Experiment Results
@@
Appendix G Full Query Expansion Experiment Results
@@
Appendix H Full Reranking Experiment Results
Dataset Linguistics R@5 R@10 R@20 R@100
@@ Appendix H Full Reranking Experiment Results
Appendix I Rewriting Prompts
I.1 Formality
@@ I.3 Politeness
Appendix J LLMs Prompts
J.1 Few-shot Prompts
@@
Appendix K Rewriting Examples
@@
Appendix L Computational Resources
Our experimental setup utilized models of varying scales: Gemma-2 (2B, 9B, 27B parameters), Llama-3.1 (8B, 70B parameters), and Qwen-2.5 (3B, 7B, 32B, 72B parameters). For retrieval, we employed ModernBERT Embed (149M parameters) and Contriever. We conducted comprehensive evaluations across 5 linguistic dimensions (4 dimensions plus 2 grammatical correctness subtypes), 4 datasets, 9 language models, and 2 retrieval systems, totaling 360 experimental configurations. Each model inference run required approximately 1.5 hours, resulting in 540 GPU hours on 16 L40S GPUs distributed across different model configurations. Retrieval evaluation required an additional 40 GPU hours. Data rewriting was performed using GPT-4o-mini, requiring 40 hours of API usage. Total computational cost comprised 620 GPU hours on L40S hardware plus commercial API usage for data preprocessing.